OpenStudy (anonymous):

(12-sqrt-4)+(7+sqrt-9)

OpenStudy (anonymous):

\[(12-\sqrt{-4})+(7+\sqrt{-9})\]

geerky42 (geerky42):

Remember, \( \sqrt{-a} = i\sqrt a\)

geerky42 (geerky42):

Do that then combine like terms.

geerky42 (geerky42):

Don't forgot to simply.

OpenStudy (anonymous):

\[2\sqrt{3}-2i+7+3i\]

OpenStudy (anonymous):

oh its 12+7=19

OpenStudy (anonymous):

19+i

geerky42 (geerky42):

Yep.